OpenStudy (anonymous):

1/r+1/s and s+r/rs evaluate both expressions by using R=200 and S=250

jhonyy9 (jhonyy9):

1/r +1/s= (s+r)/rs 1/200 +1/250 = (250 +200)/200*250 = 450/50000 = 45/5000 = 9/1000